# AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks

### v1

@@ C.4. Answering Prompts
[Steps:] Please follow these steps: 1. Analyze the Question: - Carefully read the question and its answer choices. 2. Analyze the Related TTPs: - Analyze the list of Related TTPs to understand the context of the question. 3. Compare and Evaluate: - Based on the related TTPs, evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below.

### v2

@@ C.4 Answering Prompt Templates
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the CTI report. Note that the CTI report contains key details required for your analysis, but it may not directly state the answer. Your evaluation of the answer choices is essential to arrive at the correct answer. [Steps:] Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 3. Compare and Evaluate: - Match the extracted attack sequence from the CTI report with the details in the question. - Evaluate each answer choice to determine which one aligns best with the attack sequence and any critical contextual information. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: “Final Answer: <insert answer choice here>.” [End of Steps] Following the steps above, please answer the question below using the provided CTI report.
Box 10: Prompt template for the Zero-Shot setting. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive a question about a sequence of TTPs with several answer choices. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question. [Steps:] Please follow these steps: 1. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 2. Compare and Evaluate: - Evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 3. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: “Final Answer: <insert answer choice here>.” [End of Steps] Following the steps above, please answer the question below.
Box 11: Prompt template for the RAG-empowered setting. You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A Question about a sequence of TTPs with several answer choices. 2. A list of Related TTPs that are relevant to the question. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question. [Steps:] Please follow these steps: 1. Analyze the Question: - Carefully read the question and its answer choices. 2. Analyze the Related TTPs: - Analyze the list of Related TTPs to understand the context of the question. 3. Compare and Evaluate: - Based on the related TTPs, evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: “Final Answer: <insert answer choice here>.” [End of Steps] Following the steps above, please answer the question below.